OpenStudy (anonymous):

Sharon found a couch for $1,770.00. The sales person said it was on sale for 25% off. How much did Sharon pay for the couch if she also had to pay 6% sales tax on the total? A. $469.05 B. $1,327.50 C. $1,433.70 D. $1,407.15

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

1770/4=442.5 1770-442.5=1327.5 1327.5*.06=79.65 1327.5+79.65= $1407.15 - D

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

i need help on one more if thats ok

OpenStudy (anonymous):

Tara's golf score was 23 less than twice Amber's score. If Tara's score was 93, which equation could you use to find Amber's score? Let a = Amber's score. A. 2(a - 23) = 93 B. 23 - 2a = 93 C. 2(93) - 23 = a D. 2a - 23 = 93